OpenStudy (anonymous):

Minnie has a piece of cardboard with length (3x2 − 2x + 1) inches and width (2x2 + 3x − 4) inches. Which statement best explains why the area of the cardboard demonstrates the closure property?

OpenStudy (anonymous):

Minnie has a piece of cardboard with length (3x2 − 2x + 1) inches and width (2x2 + 3x − 4) inches. Which statement best explains why the area of the cardboard demonstrates the closure property? It is equal to (6x4 + 9x3 − 12x2 + 8x − 4) square inches, which is also a polynomial like the length and width. It is equal to (6x4 + 5x3 − 16x2 + 11x − 4) square inches, which is also a polynomial like the length and width. The length multiplied by the width is equal to the width multiplied by the length. The length multiplied by the width has the same degree as the width multiplied by the length.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@uri @Preetha

OpenStudy (michele_laino):

hint: area of a rectangle is given by the subsequent formula: \[{\text{area = base }} \times {\text{ height}}\]

OpenStudy (michele_laino):

so we can write: \[\Large \begin{gathered} {\text{area = base }} \times {\text{ height = }} \hfill \\ = \left( {3{x^2} - 2x + 1} \right) \times \left( {2{x^2} + 3x - 4} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

so its gonna be 6x^4-6x^2-4 correct?

OpenStudy (michele_laino):

I got a different result

OpenStudy (anonymous):

uhmm

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \left( {3{x^2} - 2x + 1} \right) \times \left( {2{x^2} + 3x - 4} \right) = \hfill \\ = 6{x^4} + 9{x^3} - 12{x^2} - 4{x^3} - 6{x^2} - 4 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

how did you get that

OpenStudy (michele_laino):

I have applied the distributive peroperty of multiplication over addition

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so now what do i do?

OpenStudy (michele_laino):

more explanation: \[\large \begin{gathered} \left( {3{x^2} - 2x + 1} \right) \times \left( {2{x^2} + 3x - 4} \right) = \hfill \\ = 3{x^2} \times \left( {2{x^2} + 3x - 4} \right) - 2x \times \left( {2{x^2} + 3x - 4} \right) + 1 \times \left( {2{x^2} + 3x - 4} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

\[\begin{gathered} \left( {3{x^2} - 2x + 1} \right) \times \left( {2{x^2} + 3x - 4} \right) = \hfill \\ = 3{x^2} \times \left( {2{x^2} + 3x - 4} \right) - 2x \times \left( {2{x^2} + 3x - 4} \right) + 1 \times \left( {2{x^2} + 3x - 4} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

this is gonna be really drawn out

OpenStudy (anonymous):

6x^4+9x3-12x^2-4x^3-6x^2+8x+2x^2+3x-4?

OpenStudy (michele_laino):

yes! you are right! sorry I have meade an error in my previous result

OpenStudy (michele_laino):

now you have to simplify similar terms, what do you get?

OpenStudy (anonymous):

ohh crap uhmm

OpenStudy (anonymous):

can you help me on this i dont really know

OpenStudy (michele_laino):

oops.. I have made*

OpenStudy (michele_laino):

grouping similar term, we can write: \[\Large 6{x^4} + {x^3}\left( {9 - 4} \right) + {x^2}\left( { - 12 - 6 + 2} \right) + x\left( {8 + 3} \right) - 4 = ...?\]

OpenStudy (michele_laino):

\[\large 6{x^4} + {x^3}\left( {9 - 4} \right) + {x^2}\left( { - 12 - 6 + 2} \right) + x\left( {8 + 3} \right) - 4 = ...?\]

OpenStudy (anonymous):

uhmm idk

OpenStudy (michele_laino):

what is 9-4=...?

OpenStudy (anonymous):

5

OpenStudy (michele_laino):

ok! and -12-6+2=...? 8+3=...?

OpenStudy (anonymous):

-16 and 11

OpenStudy (michele_laino):

that's right! So, what is the resulting polynomial?

OpenStudy (anonymous):

what you mean

OpenStudy (michele_laino):

you have to substitute the coefficients which you have fount into this expression: \[6{x^4} + {x^3}\left( {9 - 4} \right) + {x^2}\left( { - 12 - 6 + 2} \right) + x\left( {8 + 3} \right) - 4 = ...?\]

OpenStudy (michele_laino):

found*

OpenStudy (anonymous):

so im still confused? what did we just do with the addition and subtraction?

OpenStudy (anonymous):

is the answer C?

OpenStudy (michele_laino):

the resulting polinomial is: \[\begin{gathered} \left( {3{x^2} - 2x + 1} \right) \times \left( {2{x^2} + 3x - 4} \right) = \hfill \\ \hfill \\ = 3{x^2} \times \left( {2{x^2} + 3x - 4} \right) - 2x \times \left( {2{x^2} + 3x - 4} \right) + 1 \times \left( {2{x^2} + 3x - 4} \right) = \hfill \\ \hfill \\ = 6{x^4} + {x^3}\left( {9 - 4} \right) + {x^2}\left( { - 12 - 6 + 2} \right) + x\left( {8 + 3} \right) - 4 = \hfill \\ \hfill \\ = 6{x^4} + 5{x^3} - 16{x^2} + 11x - 4 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

dang sorry i was confused but thank you so much you were a huge help!!! xD

OpenStudy (michele_laino):

it is the second option

OpenStudy (michele_laino):

please wait, I think it is the fourth option, because we got a fourth-degree polynomial, whereas the length and the width are two second-degree polynomials